OpenStudy (anonymous):

Need help I'm 10 A new clothing store had expenses of $60,000 for designing and building the shelves and counters and $140,000 for the first year's inventory. So far this year, the store has sales of $80,000. What do the sales for the rest of the year have to be for the store to break even? A. $60,000 B. $120,000 C. $140,000 D. $200,000

OpenStudy (anonymous):

A

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

B is the right answer First add expenses as negative numbers -60,000 + (-140,000) = -200,000 Since the expenses are $200,000, the store needs to sell $200,000 in clothes to break even. You know they have already sold $80,000 worth, so you can set up an equation and solve for the missing dollar amount they need to sell: 80,000 + s = 200,000 s = 200,000 - 80,000 s = 120,000 The store needs to sell $120,000 more in clothes to break even.